# Lost in Tracking Translation: A Comprehensive Analysis of Visual SLAM in Human-Centered XR and IoT Ecosystems

### v1

@@ 5.1. Input Profiling
The HoloSet dataset introduces human-centered sequences with both indoor and outdoor settings. These sequences feature a human wearing an XR headset, resulting in slower and more unpredictable motion than the drone or car-based sequences in EuRoC and KITTI. For the Campus Center sequences, motion is slow and set indoors with medium lighting. Here, systems can focus on identifying and tracking key features in cluttered scenes with human movement and stairs while adjusting to the slower motion. In contrast, suburb jog sequences involve fast motion in bright, outdoor environments with trees, parked cars, and roads. Profiling these sequences shows the need for systems to anticipate rapid changes in scene structure and motion while handling outdoor lighting challenges. Like KITTI, environmental elements like trees introduce occlusion challenges, and tracking algorithms can prioritize high-texture regions while compensating for motion blur and abrupt human movements.
By tailoring tracking methods to the specific conditions of each sequence—whether it involves fast motion, complex geometry, or challenging lighting—algorithms can dynamically adjust their processing techniques. For example:
@@
Table 8. Profiling Lighting and Texture on EuRoC.
Sequence MH01 MH02 MH03 MH04 MH05 V101 V102 V103 V201 V202 V203
@@ 5.3. Output Evaluation
Cross-Domain Adaptation. By analyzing output errors across different domains (e.g., indoor to outdoor transitions in EuRoC vs. KITTI or the human-centric scenes in HoloSet), systems can gain insights into how to improve domain adaptability. Output evaluation helps to pinpoint weaknesses that may not be evident during training but become apparent in real-world, cross-domain deployments.
By systematically evaluating the outputs, SLAM systems can continuously improve their tracking accuracy, refining their methods based on real-world data and improving their resilience across different environments.

### v2

@@ 5.1. Input Profiling
Table 11. Profiling Lighting and Texture on EuRoC.
Sequence MH01 MH02 MH03 MH04 MH05 V101 V102 V103 V201 V202 V203
The HoloSet dataset introduces human-centered sequences with both indoor and outdoor settings. These sequences feature a human wearing an XR headset, resulting in slower and more unpredictable motion than the drone or car-based sequences in EuRoC and KITTI. For the Campus Center sequences, motion is slow and set indoors with medium lighting. Here, systems can focus on identifying and tracking key features in cluttered scenes with human movement and stairs while adjusting to the slower motion. In contrast, suburb jog sequences involve fast motion in bright, outdoor environments with trees, parked cars, and roads. Profiling these sequences shows the need for systems to anticipate rapid changes in scene structure and motion while handling outdoor lighting challenges. Like KITTI, environmental elements like trees introduce occlusion challenges, and tracking algorithms can prioritize high-texture regions while compensating for motion blur and abrupt human movements. By tailoring tracking methods to the specific conditions of each sequence—whether it involves fast motion, complex geometry, or challenging lighting—algorithms can dynamically adjust their processing techniques. For example:
@@
5.2. Sensor-Specific Observations
Tracking performance is shaped not only by algorithmic design and scene characteristics but also by the properties and limitations of the sensors used for visual tracking. In our evaluation, we observed that several sensor factors interact with environmental and locomotion challenges to affect SLAM and VO accuracy.
Shutter effects. Rolling shutter distortion is a significant source of error in HoloSet sequences, where fast head motions introduce geometric artifacts. Algorithms that lack explicit rolling shutter compensation struggle to maintain accuracy under these conditions, with many SLAM methods failing entirely.
Resolution and field-of-view (FoV). Narrow FoV and limited resolution in the HoloLens-based HoloSet dataset reduce feature coverage, particularly in texture-sparse indoor scenes and during rapid rotations. A narrow FoV also exacerbates sensitivity to rotational motion, leading to larger drift. This affects both traditional feature-based methods and deep learning methods.
Frame rate and synchronization. High frame rate sequences in HoloSet (60 FPS) impose additional computational demands and require accurate temporal synchronization. Misalignment between camera and IMU streams, or dropped frames under processing load, introduce drift and instability. Some methods, such as VINS-Fusion, show degraded performance in high-motion HoloSet sequences compared to controlled EuRoC runs. In contrast, EuRoC’s global shutter stereo cameras with well-synchronized IMU yield more stable performance.
Dynamic range and lighting sensitivity. The strong correlation of ATE with lighting variations (Table 8) highlights the influence of sensor dynamic range and exposure control. Methods that rely on photometric consistency are particularly vulnerable to lighting changes and overexposure. KITTI’s outdoor sequences show pronounced variability in ATE under varying sun angles and shadows.
Sensor noise and calibration. Sensor noise, including image noise and IMU bias, also impacts tracking accuracy. HoloSet’s IMU noise and drift, in combination with rolling shutter effects, challenge visual-inertial methods. In KITTI, well-calibrated stereo rigs enable robust tracking even under vehicle vibrations. EuRoC similarly benefits from high-quality calibration.
Lens distortion and optical artifacts. Optical distortion, particularly at the edges of wide-angle lenses, can introduce tracking instability if not properly modeled. In HoloSet and EuRoC, lens distortion interacts with narrow FoV and rolling shutter effects, degrading feature tracking in peripheral image regions.
Sensor limitations compound tracking challenges in complex environments across all datasets. In HoloSet, rolling shutter distortion, narrow FoV, high frame rates, and sensor noise significantly degrade tracking accuracy. In KITTI, the global shutter mitigates geometric distortion, but dynamic lighting and vehicle vibration remain key challenges. EuRoC’s well-calibrated, synchronized sensor suite yields the most consistent performance, though sensitivity to motion blur and lighting remains. Robust SLAM and VO methods must explicitly account for sensor-specific effects—particularly when targeting deployment on heterogeneous hardware platforms as found in XR and IoT ecosystems.
@@ 5.4. Output Evaluation
Cross-Domain Adaptation. By analyzing output errors across different domains (e.g., indoor-to-outdoor transitions in EuRoC vs. KITTI or human-centric scenes in HoloSet), systems can gain actionable insights into how to improve domain adaptability. Output evaluation helps pinpoint weaknesses that may not be evident during training but become apparent in real-world, cross-domain deployments. A promising direction to improve robustness is the use of cross-domain adaptation techniques that enable models to generalize better to unseen environments. Recent approaches such as self-supervised deep visual odometry with online adaptation(Li et al., 2020) and adaptive learning for hybrid visual odometry(Liu et al., 2024) explicitly address this challenge, showing how online adaptation and domain-invariant representations can improve performance in diverse scenes.
By systematically evaluating outputs and integrating such adaptation strategies, SLAM systems can continuously refine their tracking accuracy, becoming more resilient across varying environments.